OpenStudy (anonymous):

Solve for x 3x+3y=15 3x+2y=13 Medal and Fan!!

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

\[3x+3y=15\] \[3(x+y)=15\] \[x+y=5\] ??

OpenStudy (anonymous):

????

OpenStudy (anonymous):

i dont get it!!!!!

hartnn (hartnn):

i'd say, try subtracting the 2nd equation from the 1st equation, what u get ?

OpenStudy (anonymous):

Okay so i at least know that you have to solve for x in the first equation then substitute x for the second equation. But i dont know how to do that!!

hartnn (hartnn):

thats substitution method what i suggested was elimination method if you are more comfortable with substitution method, you know you need to solve for x have you tried, which step are u stuck ?

OpenStudy (anonymous):

i just dont know how to solve for x in the first equation

hartnn (hartnn):

start by factoring out 3 from left side 3x + 3y = 3 (...?..)

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

would i subtract 3 from both sides???

OpenStudy (anonymous):

im confused!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (phi):

When you see 3x+3y=15 3x+2y=13 you should notice that 3x - 3x is 0 (and the x's go away) if you subtract the top equation minus the bottom equation, you will get a new (improved!) equation with only a y. solve for y, then use that value in either of the original equations to find x.

OpenStudy (anonymous):

(3x+3y=15) multiply by 2 (3x+2y=13) multiply by -3 result: 6x+6y=30 -9x-6y=-39 add the two equation using elimination method -3x=-9 x=3

OpenStudy (phi):

subract an equation means: multiply every term in the bottom equation by -1, then add term by term to the top equation

OpenStudy (anonymous):

mine is a bit more complex than @phi... :D

OpenStudy (anonymous):

y=2?

OpenStudy (anonymous):

correct

OpenStudy (phi):

Now use y=2 to find x

OpenStudy (anonymous):

just substitute it to any of the 2 given eqation and you get x. :)

OpenStudy (anonymous):

x=3?